Hot Wheels Cycle Shop bought 16 helmets at $30 each.
Some were sold for $60 each.
The rest were sold at a clearance sale price of $45 each.
The profit for Hot Wheels was $375.
How many helmets were sold at the sale price?
The word problem says "Cost and Money Problems". If that helps at all.

Answer by Prithwis(166)

Total Cost of helmets = 16*30 = $480;
Total Profit = $375;
Total Sell Price = Total Cost + Total Profit = 480+375 = $855;
Assume x helmets were sold for $60 each; 16-x helmets were sold on clearance;
Total money reveived by selling x helmets for $60 each = 60x;
Remaining money = Total Sell Price - Total price of x helmets = 855-60x
Number of helmets sold on clearance*Clearance Price of helmet = Remaining money
So, (16-x)(45) = 855-60x
=> (16-x)(5*9) = 5(171-12x)
=> 9(16-x) = 171-12x
=> 144-9x = 171-12x
=> 3x = 171-144 = 27
=> x = 9;
Answer - Number of helmets sold at the sale price was 9